The Ship Corp. has paid annual dividends of $0.48, $0.60, and $0.62 a share over the past
three years, respectively. The company now predicts that it will maintain a constant
dividend since its business has leveled off and sales are expected to remain relatively
constant. Given the lack of future growth, at what price will you only buy this stock if you
can earn at least a 14 percent rate of return?
1
Expert's answer
2016-02-21T00:01:09-0500
Dividends of $0.48, $0.60, and $0.62 a share over the past three years, 14 percent rate of return.
If the last year dividents were $0.62 per share, the expected rate of return is 14% and there will be no growth in divedents in the future, then the price of the stock will be: P = D/(k - g) = 0.62/(0.14 - 0) = $4.43 per share.